# Shen2018S18

### v1

@@ 4 Spectral energy distributions and luminous star masses
We fit the SEDs using MINESweeper (Cargile et al., 2020), a code designed for joint modeling of stellar photometry and spectra. We only use the photometric modeling capabilities of the code but place a prior on the present-day metallicity from spectroscopy. The free parameters of the fit are the initial mass and metallicity of each star, its age (as parameterized by the “equivalent evolutionary point”), the parallax, and the foreground extinction. We place priors on the parallax from the Gaia 12-parameter binary solution and set an age upper limit of 13 Gyr. We also use a metallicity prior from the high-resolution spectra (Table 2), and an extinction prior from 3D dust maps. We again use the Green et al. (2019) map for sources with declination $\delta>-30$ deg, and the Lallement et al. (2022) map for sources farther south.
@@ 6.2 Period-eccentricity relation
Although the eccentricities of the WD+MS binary candidates in Figure 12 are much lower on average than those of ordinary MS+MS binaries at the same periods, the large majority of them are not consistent with zero, and are much higher than expected for binaries that have gone through long periods of stable mass transfer (Phinney, 1992; Lorimer, 2008). This may be a consequence of these systems having first initiated mass transfer on the asymptotic giant branch rather than on the first giant branch and/or having gone through common envelope evolution rather than stable mass transfer.
@@ 6.3 Galactic orbits
The results are shown in Figure 13. We compare the NS candidates in our sample to all binaries with AstroSpectroSB1 solutions (left panels) and to those with $d=0.4-1.0$ kpc (right panels; these have a similar distance distribution to the NS candidates). Dashed lines centered on $V_{\phi}=225\,\rm km\,s^{-1}$ show approximate boundaries of the Galactic thin disk (total velocity $<70\,\rm km\,s^{-1}$), thick disk ($70-180\,\rm km\,s^{-1}$), and halo ($>180\,\rm km\,s^{-1}$) (e.g. Bensby et al., 2014). Three objects in our sample – J1739+4502, J1432-1021, and J0152-2049 – are on halo orbits, with total velocities of 370, 350, and 290 $\rm km\,s^{-1}$ with respect to the local standard of rest. These objects all have metallicities $\rm[Fe/H]<-1.2$, implying that their high velocities are mainly the result of membership to an old stellar population, not natal kicks. On the other hand, the 6 systems in the “thick disk” region of the Toomre diagram have metallicities close to solar. This suggests that natal kicks may have played a role in kinematically heating their orbits.
@@ 6.4 Lithium enhancement
Our inferred $\rm A(Li)$ are higher than the predicted yields of super-AGB stars with $M\lesssim 8\,M_{\odot}$, but compatible with predictions for $8\,M_{\odot}$ models (see Ventura & D’Antona 2010, who note that these yields depend sensitively on the assumed winds). This could support a scenario where the companions are NSs formed through electron-capture supernovae, or ultramassive WDs. However, this scenario is difficult to reconcile with the fact that one of the binaries in question has $M_{2}=1.90\pm 0.03\,M_{\odot}$, whereas electron capture SNe are expected to produce low-mass NSs.
@@ 6.5 Constraints on kicks and mass loss
Models predict that dynamical channels are less efficient for forming wide NS binaries than for BH binaries (Tanikawa et al., 2024). Therefore, we assume the systems in our sample formed from primordial binaries and use their current orbits to place limits on their pre-supernova masses and natal kicks.
@@
Curiously, the most massive NS in the sample, Gaia NS1 (J1432-1021), has the lowest-eccentricity orbit, requiring the lowest kicks. This appears inconsistent with SNe simulations, which predict high-mass NSs to receive the strongest kicks (Burrows et al., 2023). It is possible that Gaia NS1 formed through a different channel than the other objects in the sample, but larger sample is required to investigate this possibility quantitatively.
@@ 6.6 Future evolution to symbiotic X-ray binaries
At least a few neutron stars with solar-mass red giant companions are known to exist in symbiotic X-ray binaries (e.g. Hinkle et al., 2006, 2019). One system that is likely representative of our candidates’ future evolution is GX 1+4 (Davidsen et al., 1977; Hinkle et al., 2006), which contains a $\lesssim 1\,M_{\odot}$ giant in a 1161-day orbit at a distance of $\approx 4$ kpc. Another similar system is the symbiotic X-ray binary IGR J16194-2810 (Masetti et al., 2007), which contains a $\sim 1\,M_{\odot}$ red giant at a distance of $\approx 2.1$ kpc in a $\sim 200$-day orbit with a NS companion (Nagarajan et al. 2024, in prep). Given their X-ray properties, these systems undoubtedly contain NSs. This provides some support for the interpretation that the dark objects in our binaries are also NSs, despite the uncertainties associated with forming NS+MS binaries in such wide orbits.
@@ 7 Conclusions
    Quality of the Gaia orbital solutions: RVs for all the objects in our final sample are in good agreement with predictions of the Gaia orbital solutions (Figure 5), and joint fitting of RVs and astrometry leads to well-constrained orbits (Figure 6). For most objects, constraints from the Gaia-only and Gaia+RVs solutions are consistent within $1\sigma$ (Figure 7). On average, companion masses constrained by joint RV+astrometry fits are lower than the pure-astrometry estimates (Figure 8); this is an expected consequence of selecting candidates from the upper tail of the companion mass distribution. The objects all have unproblematic Gaia quality flags and astrometric uncertainties typical of their apparent magnitudes (Figure 10). Most of the orbits are predicted to have been sampled uniformly in phase by Gaia (Figure 9).
@@ acknowledgments
We thank Josh Simon, Casey Lam, Kyle Kremer, Jim Fuller, and Thomas Tauris for useful discussions. We are grateful to Yuri Beletsky, Sam Kim, Angela Hempel, Régis Lachaume, Gil Esquerdo, Perry Berlind, and Mike Calkins for observing help. This research was supported by NSF grant AST-2307232. HWR acknowledges support from the European Research Council for the ERC Advanced Grant [101054731]. This research benefited from discussions in the ZTF Theory Network, funded in part by the Gordon and Betty Moore Foundation through Grant GBMF5076, and from collaboration at the “Renaissance of Stellar Black-Hole Detections in The Local Group” workshop hosted at the Lorentz Center in June, 2023.
@@
Appendix A Rejection of spurious solutions
@@ Appendix C All fits
Orbital period $P_{\rm orb}$ [days] $480.9\pm 0.6$
Semi-major axis $a$ [au] $1.624\pm 0.011$
Photocenter semi-major axis $a_{0}$ [mas] $1.441\pm 0.020$
Eccentricity $e$ $0.432\pm 0.017$
Inclination $i$ [deg] $136.4\pm 1.0$
Periastron time $T_{p}$ [JD-2457389] $138.1\pm 1.5$
Ascending node angle $\Omega$ [deg] $229.4\pm 1.4$
Argument of periastron $\omega$ [deg] $214.1\pm 1.2$
Luminous star mass $M_{\star}$ [$M_{\odot}$] $1.029\pm 0.020$
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.44\pm 0.04$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $16.4\pm 0.5$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.160\pm 0.013$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $18.1\pm 0.4$
Parallax $\varpi$ [mas] $1.520\pm 0.014$
@@
Orbital period $P_{\rm orb}$ [days] $631.80\pm 0.23$
Semi-major axis $a$ [au] $1.973\pm 0.025$
Photocenter semi-major axis $a_{0}$ [mas] $1.860\pm 0.018$
Eccentricity $e$ $0.553\pm 0.004$
@@
Argument of periastron $\omega$ [deg] $180.36\pm 0.15$
@@
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $18.68\pm 0.30$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.246\pm 0.010$
@@
Orbital period $P_{\rm orb}$ [days] $632.76\pm 0.22$
@@
Photocenter semi-major axis $a_{0}$ [mas] $1.965\pm 0.012$
Eccentricity $e$ $0.2635\pm 0.0013$
Inclination $i$ [deg] $105.6\pm 0.5$
@@
Ascending node angle $\Omega$ [deg] $132.7\pm 0.5$
Argument of periastron $\omega$ [deg] $55.67\pm 0.16$
@@
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.261\pm 0.032$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $18.611\pm 0.019$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.3794\pm 0.0012$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $-69.667\pm 0.017$
Parallax $\varpi$ [mas] $1.812\pm 0.012$
@@
Semi-major axis $a$ [au] $1.912\pm 0.018$
Photocenter semi-major axis $a_{0}$ [mas] $1.376\pm 0.017$
Eccentricity $e$ $0.6770\pm 0.0019$
Inclination $i$ [deg] $144.6\pm 0.5$
Periastron time $T_{p}$ [JD-2457389] $179.6\pm 2.3$
Ascending node angle $\Omega$ [deg] $70.6\pm 1.5$
Argument of periastron $\omega$ [deg] $334.43\pm 0.22$
Luminous star mass $M_{\star}$ [$M_{\odot}$] $0.778\pm 0.030$
@@
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $15.910\pm 0.034$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.1094\pm 0.0007$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $-264.214\pm 0.027$
@@
Orbital period $P_{\rm orb}$ [days] $814.2\pm 3.4$
Semi-major axis $a$ [au] $2.370\pm 0.028$
Photocenter semi-major axis $a_{0}$ [mas] $1.338\pm 0.025$
Eccentricity $e$ $0.381\pm 0.016$
Inclination $i$ [deg] $106.7\pm 1.0$
Periastron time $T_{p}$ [JD-2457389] $-56.1\pm 3.3$
Ascending node angle $\Omega$ [deg] $160.9\pm 1.4$
Argument of periastron $\omega$ [deg] $284.6\pm 1.0$
Luminous star mass $M_{\star}$ [$M_{\odot}$] $1.00\pm 0.05$
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.68\pm 0.07$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $20.6\pm 0.7$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.58\pm 0.04$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $60.1\pm 0.9$
Parallax $\varpi$ [mas] $0.899\pm 0.012$
@@
Orbital period $P_{\rm orb}$ [days] $889.8\pm 1.2$
Semi-major axis $a$ [au] $2.407\pm 0.030$
Photocenter semi-major axis $a_{0}$ [mas] $5.93\pm 0.04$
Eccentricity $e$ $0.583\pm 0.010$
Inclination $i$ [deg] $125.33\pm 0.26$
Periastron time $T_{p}$ [JD-2457389] $-392.4\pm 0.8$
@@
Argument of periastron $\omega$ [deg] $243.2\pm 0.4$
@@
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.40\pm 0.04$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $17.59\pm 0.28$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.269\pm 0.007$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $-43.23\pm 0.17$
Parallax $\varpi$ [mas] $4.135\pm 0.017$
@@
Orbital period $P_{\rm orb}$ [days] $1027\pm 4$
Semi-major axis $a$ [au] $2.775\pm 0.020$
Photocenter semi-major axis $a_{0}$ [mas] $2.70\pm 0.04$
Eccentricity $e$ $0.686\pm 0.013$

### v2

@@ 3.9.4 Metallicity comparisons
We also investigated the consistency of metallicities measured from high-resolution spectra by different pipelines. For sources with more than one high resolution spectrum, we found a median absolute deviation of 0.08 dex between measurements, which is comparable to the reported uncertainties.
@@ 4 Spectral energy distributions and luminous star masses
We fit the SEDs using MINESweeper (Cargile et al., 2020), a code designed for joint modeling of stellar photometry and spectra. We only use the photometric modeling capabilities of the code but place a prior on the present-day metallicity from spectroscopy. The free parameters of the fit are the initial mass and metallicity of each star, its age (as parameterized by the “equivalent evolutionary point”), the parallax, and the foreground extinction. For each call to the likelihood function, the mass, metallicity, and equivalent evolutionary point are converted to a radius and effective temperature using MIST isochrones (Choi et al., 2016), and are then used to predict a model SED that is compared to the data. We place priors on the parallax from the Gaia 12-parameter binary solution and set an age upper limit of 13 Gyr. We also use a metallicity prior from the high-resolution spectra (Table 2), and an extinction prior from 3D dust maps. We again use the Green et al. (2019) map for sources with declination $\delta>-30$ deg, and the Lallement et al. (2022) map for sources farther south.
@@ 6.2 Period-eccentricity relation
Although the eccentricities of the WD+MS binary candidates in Figure 12 are much lower on average than those of ordinary MS+MS binaries at the same periods, the large majority of them are not consistent with zero, and are much higher than expected for binaries that have gone through long periods of stable mass transfer (Phinney, 1992; Lorimer, 2008). This may be a consequence of these systems having first initiated mass transfer on the asymptotic giant branch rather than on the first giant branch and/or having gone through common envelope evolution rather than stable mass transfer. Similar eccentricities are observed in other populations of WD+MS binaries at these periods, such as blue stragglers and barium stars (Mathieu & Geller, 2009; Jorissen et al., 2019; Escorza et al., 2019).
@@ 6.3 Galactic orbits
The results are shown in Figure 13. We compare the NS candidates in our sample to all binaries with AstroSpectroSB1 solutions (left panels) and to those with $d=0.4-1.0$ kpc (right panels; these have a similar distance distribution to the NS candidates). Dashed lines centered on $V_{\phi}=225\,\rm km\,s^{-1}$ show approximate boundaries of the Galactic thin disk (total velocity $<70\,\rm km\,s^{-1}$), thick disk ($70-180\,\rm km\,s^{-1}$), and halo ($>180\,\rm km\,s^{-1}$) (e.g. Bensby et al., 2014). Three objects in our sample – J1739+4502, J1432-1021, and J0152-2049 – are on halo orbits, with total velocities of 370, 350, and 290 $\rm km\,s^{-1}$ with respect to the local standard of rest. These objects all have metallicities $\rm[Fe/H]<-1.2$, implying that their high velocities are mainly the result of membership to an old stellar population, not natal kicks. On the other hand, the 6 systems in the “thick disk” region of the Toomre diagram have metallicities close to solar. This suggests that natal kicks may have played a role in kinematically heating their orbits. The median and middle 68% range of midplane distances, $|z|$, are respectively 0.37 kpc and $(0.11-0.55)$ kpc. This makes the objects in our sample kinematically cold compared to NS LMXBs, which typically have $|z|\gtrsim 1\,{\rm kpc}$ (van Paradijs & White, 1995; Jonker & Nelemans, 2004). This is not surprising since only NSs formed with weak kicks will remain bound in the wide orbits to which astrometry is sensitive.
@@ 6.4 Lithium enhancement
Our inferred $\rm A(Li)$ are higher than the predicted yields of super-AGB stars with $M\lesssim 8\,M_{\odot}$, but compatible with predictions for $8\,M_{\odot}$ models (see Ventura & D’Antona 2010, who note that these yields depend sensitively on the assumed winds). This could support a scenario where the companions are NSs formed through electron-capture SNe, or ultramassive WDs. However, this scenario is difficult to reconcile with the fact that one of the binaries in question has $M_{2}=1.90\pm 0.03\,M_{\odot}$, whereas electron capture SNe are expected to produce low-mass NSs.
@@ 6.5 Constraints on kicks and mass loss
Models predict that dynamical channels are less efficient for forming wide NS binaries than for BH binaries (Tanikawa et al., 2024). Therefore, we assume the systems in our sample formed from primordial binaries and use their current orbits to place limits on their pre-SN masses and natal kicks.
@@
Curiously, the most massive NS in the sample, Gaia NS1 (J1432-1021), has the lowest-eccentricity orbit, requiring the lowest kicks. This appears inconsistent with SNe simulations, which predict high-mass NSs to receive the strongest kicks (Müller et al., 2019; Burrows et al., 2023). It is possible that Gaia NS1 formed through a different channel than the other objects in the sample, but larger sample is required to investigate this possibility quantitatively.
@@ 6.6 Future evolution to symbiotic X-ray binaries
At least a few neutron stars with solar-mass red giant companions are known to exist in symbiotic X-ray binaries (e.g. Hinkle et al., 2006, 2019). One system that is likely representative of our candidates’ future evolution is GX 1+4 (Davidsen et al., 1977; Hinkle et al., 2006), which contains a $\lesssim 1\,M_{\odot}$ giant in a 1161-day orbit at a distance of $\approx 4$ kpc. Another similar system is the symbiotic X-ray binary IGR J16194-2810 (Masetti et al., 2007), which contains a $\sim 1\,M_{\odot}$ red giant at a distance of $\approx 2.1$ kpc in a 193-day orbit with a NS companion (Nagarajan et al., 2024; Hinkle et al., 2024). Given their X-ray properties, these systems undoubtedly contain NSs. This provides some support for the interpretation that the dark objects in our binaries are also NSs, despite the uncertainties associated with forming NS+MS binaries in such wide orbits.
@@ 7 Conclusions
    Quality of the Gaia orbital solutions: RVs for all the objects in our final sample are in good agreement with predictions of the Gaia orbital solutions (Figure 5), and joint fitting of RVs and astrometry leads to well-constrained orbits (Figure 6). For most objects, constraints from the Gaia-only and Gaia+RVs solutions are consistent within $1\sigma$ (Figure 7). However, some initial candidates were removed from the sample after RV follow-up revealed tensions with the Gaia solution (Figure 19 and Table 3). Most but not all of these objects have large goodness_of_fit, indicative of a problematic astrometric solution. On average, companion masses constrained by joint RV+astrometry fits are lower than the pure-astrometry estimates (Figure 8); this is an expected consequence of selecting candidates from the upper tail of the companion mass distribution. The objects all have unproblematic Gaia quality flags and astrometric uncertainties typical of their apparent magnitudes (Figure 10). Most of the orbits are predicted to have been sampled uniformly in phase by Gaia (Figure 9).
@@ acknowledgments
We thank the referee for a constructive report, and Josh Simon, Casey Lam, Kyle Kremer, Jim Fuller, and Thomas Tauris for useful discussions. We are grateful to Yuri Beletsky, Sam Kim, Angela Hempel, Régis Lachaume, Gil Esquerdo, Perry Berlind, and Mike Calkins for observing help. This research was supported by NSF grant AST-2307232. HWR acknowledges support from the European Research Council for the ERC Advanced Grant [101054731]. This research benefited from discussions in the ZTF Theory Network, funded in part by the Gordon and Betty Moore Foundation through Grant GBMF5076, and from collaboration at the “Renaissance of Stellar Black-Hole Detections in The Local Group” workshop hosted at the Lorentz Center in June, 2023.
@@
Appendix A Rejection of spurious solutions and summary of all follow-up
@@
In Table 3, we provide a summary of our follow-up of all NS candidates from Shahaf et al. (2023b) and Andrews et al. (2022). For the Shahaf et al. (2023b) sample, we list candidates for which they inferred $M_{2}>1.25\,M_{\odot}$. The Andrews et al. (2022) sample only includes targets for which they infer $M_{2}>1.4\,M_{\odot}$. A majority of the candidates ruled out by RV follow-up have significantly higher goodness_of_fit than typical sources of the same apparent magnitude (Figure 10). Several candidates with incomplete RV follow-up, particularly those with $G<15$ that are inaccessible with FEROS and TRES, remain to be studied.
Gaia DR3 Source ID Reference $P_{\rm orb}$ $G$ GoF status
[d] [mag]
1522897482203494784 S23 $45.516\pm 0.005$ 11.05 1.85 likely ultramassive WD (Yamaguchi et al., 2024)
3509370326763016704 S23 $109.4\pm 0.1$ 12.47 5.22 ruled out by RV follow-up
6281177228434199296 S23 $153.9\pm 0.4$ 11.26 8.05 ruled out by RV follow-up
4482912934572480384 S23 $182.4\pm 0.4$ 12.35 20.26 ruled out by RV follow-up
2995961897685517312 S23 $189.8\pm 0.3$ 13.00 0.41 one of our candidates
2080945469200565248 S23 $196.6\pm 0.5$ 13.29 -1.25 RV follow-up implies $e<0.001$. Likely ultramassive WD.
2032579979951732736 S23 $215.4\pm 0.4$ 14.21 -1.04 high $v\sin i$ prevented robust RV measurements
6481502062263141504 S23 $229.7\pm 0.5$ 13.58 2.11 one of our candidates
5820382041374661888 S23 $311.1\pm 1.4$ 14.19 -2.49 one of our candidates
1525829295599805184 A22 $328.2\pm 0.8$ 16.18 2.25 fainter than $G=15$
1871419337958702720 S23 $479.3\pm 1.0$ 13.70 -0.71 one of our candidates
5530442371304582912 S23 $498.8\pm 2.8$ 14.60 -2.04 one of our candidates
3263804373319076480 S23 $510.7\pm 4.7$ 12.67 5.56 ruled out by RV follow-up
6601396177408279040 S23 $533.5\pm 2.0$ 14.07 13.80 not followed-up due to poor GoF
5136025521527939072 S23 $536.9\pm 1.2$ 12.05 4.68 one of our candidates
2912474227443068544 S23 $541.5\pm 3.9$ 13.81 24.83 ruled out by RV follow-up
4271998639836225920 A22 $545.3\pm 1.1$ 15.62 0.12 fainter than $G=15$
4922744974687373440 S23 $562.4\pm 2.0$ 14.48 1.23 one of our candidates
6001459821083925120 A22 $563.8\pm 6.7$ 13.60 -1.54 ruled out by RV follow-up
3869650535947137920 S23 $570.0\pm 5.3$ 12.94 5.46 ruled out by RV follow-up (Figure 19)
1434445448240677376 S23 $572.4\pm 1.8$ 13.65 -2.97 one of our candidates
6802561484797464832 S23 $574.8\pm 6.2$ 12.88 11.01 not followed-up due to poor GoF
2196619383835483648 S23 $582.2\pm 5.0$ 15.71 1.53 fainter than $G=15$
1695294922548180224 A22, S23 $601.2\pm 2.9$ 13.12 -2.74 RV follow-up not yet completed
4744087975990080896 A22, S23 $631.3\pm 5.5$ 17.07 -0.10 fainter than $G=15$
1694708646628402048 S23 $632.0\pm 2.8$ 13.20 0.77 one of our candidates
3494029910469026432 S23 $632.5\pm 1.6$ 12.66 2.64 one of our candidates
4637171465304969216 S23 $639.2\pm 4.4$ 14.01 0.77 one of our candidates
5580526947012630912 A22, S23 $654.3\pm 4.9$ 13.36 0.66 one of our candidates
1350295047363872512 A22, S23 $657.2\pm 4.3$ 13.52 3.53 one of our candidates
6593763230249162112 A22 $679.9\pm 2.8$ 13.54 1.67 ruled out by RV follow-up (Figure 19)
4240540718818313984 S23 $691.2\pm 2.0$ 14.61 0.41 RV follow-up not yet completed
2426116249713980416 S23 $711.7\pm 21.5$ 13.02 -0.30 one of our candidates
4578398926673187328 S23 $728.1\pm 22.9$ 15.02 0.05 fainter than $G=15$
6328149636482597888 A22, S23 $736.0\pm 11.6$ 13.34 0.74 one of our candidates
2885872059004028800 S23 $802.2\pm 28.1$ 15.31 1.93 fainter than $G=15$
6037767138131854592 S23 $804.7\pm 29.2$ 14.28 -0.91 RV follow-up not yet completed
5590962927271507712 A22, S23 $817.9\pm 5.1$ 15.88 1.86 fainter than $G=15$
1058875159778407808 A22, S23 $835.7\pm 29.3$ 14.52 -0.01 one of our candidates
5446310318525312768 S23 $866.6\pm 10.6$ 10.37 4.27 RV follow-up not yet completed
6092954989675820416 S23 $883.1\pm 101.5$ 13.67 0.15 ruled out by RV follow-up
3649963989549165440 A22 $892.5\pm 60.2$ 14.30 0.52 sdB + NS or WD binary (Geier et al., 2023)
1801110822095134848 S23 $893.6\pm 2.9$ 12.19 3.58 one of our candidates
4638295715945158144 S23 $915.4\pm 5.8$ 12.93 7.68 RV follow-up not yet completed
2397135910639986304 A22, S23 $916.0\pm 19.1$ 13.35 2.10 one of our candidates
809741149368202752 A22, S23 $922.4\pm 50.5$ 14.91 1.19 RV follow-up not yet completed
1581117310088807552 A22, S23 $927.3\pm 5.7$ 14.51 2.88 one of our candidates
1749013354127453696 A22 $932.1\pm 77.5$ 14.49 -0.34 ruled out by RV follow-up
6588211521163024640 S23 $943.3\pm 44.9$ 14.19 4.92 ruled out by RV follow-up
5681911574178198400 A22, S23 $943.8\pm 31.1$ 15.64 -0.41 fainter than $G=15$
5693240254808387584 S23 $944.4\pm 85.1$ 16.40 -1.39 fainter than $G=15$
1028887114002082432 S23 $994.4\pm 40.9$ 13.59 -0.42 one of our candidates
747174436620510976 A22, S23 $999.4\pm 26.5$ 13.99 -0.96 ruled out by RV follow-up (Figure 19)
5593444799901901696 A22 $1038.8\pm 146.0$ 14.42 0.25 high $v\sin i$ prevented robust RV measurements
465093354131112960 this work $1046.3\pm 45.5$ 13.1 -0.86 one of our candidates
4314242838679237120 A22 $1146.0\pm 191.1$ 17.02 -0.78 fainter than $G=15$
1947292821452944896 A22 $1245.6\pm 163.5$ 15.94 -0.49 fainter than $G=15$
5847919241396757888 A22 $1254.3\pm 145.2$ 16.90 1.51 fainter than $G=15$
1144019690966028928 A22 $1401.9\pm 61.7$ 13.57 -1.65 RV follow-up not yet completed
1854241667792418304 A22 $1430.3\pm 33.2$ 14.87 1.38 RV follow-up not yet completed
Table 3: All the NS candidates from Andrews et al. (2022, A22) and Shahaf et al. (2023b, S23). We list the orbital period according to the Gaia astrometric solution, the apparent magnitude and goodness_of_fit, and the status of our follow-up. Candidates for which the status is colored in green are modeled in detail in this work. Those for which status is colored in red are disfavored by RV follow-up. Blue text indicates that detailed follow-up is presented elsewhere, and black text indicates that insufficient RVs have been obtained for a verdict to be reached.
@@ Appendix C All fits
Orbital period $P_{\rm orb}$ [days] $481.04\pm 0.26$
Semi-major axis $a$ [au] $1.632\pm 0.010$
Photocenter semi-major axis $a_{0}$ [mas] $1.459\pm 0.017$
Eccentricity $e$ $0.448\pm 0.009$
Inclination $i$ [deg] $135.7\pm 0.8$
Periastron time $T_{p}$ [JD-2457389] $138.2\pm 1.3$
Ascending node angle $\Omega$ [deg] $230.0\pm 1.1$
Argument of periastron $\omega$ [deg] $214.9\pm 0.9$
Luminous star mass $M_{\star}$ [$M_{\odot}$] $1.033\pm 0.03$
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.473\pm 0.034$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $16.9\pm 0.5$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.173\pm 0.011$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $18.08\pm 0.09$
Parallax $\varpi$ [mas] $1.521\pm 0.013$
Goodness of fit $\chi_{{\rm RVs}}^{2}/N_{{\rm RVs}}$ $0.5$
@@
Orbital period $P_{\rm orb}$ [days] $631.81\pm 0.22$
Semi-major axis $a$ [au] $1.973\pm 0.023$
Photocenter semi-major axis $a_{0}$ [mas] $1.859\pm 0.018$
Eccentricity $e$ $0.552\pm 0.004$
@@
Argument of periastron $\omega$ [deg] $180.38\pm 0.16$
@@
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $18.66\pm 0.28$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.246\pm 0.009$
@@
Orbital period $P_{\rm orb}$ [days] $632.65\pm 0.21$
@@
Photocenter semi-major axis $a_{0}$ [mas] $1.961\pm 0.010$
Eccentricity $e$ $0.2668\pm 0.0010$
Inclination $i$ [deg] $105.7\pm 0.4$
@@
Ascending node angle $\Omega$ [deg] $132.5\pm 0.6$
Argument of periastron $\omega$ [deg] $55.36\pm 0.14$
@@
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.258\pm 0.032$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $18.584\pm 0.018$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.3767\pm 0.0010$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $-69.617\pm 0.012$
Parallax $\varpi$ [mas] $1.812\pm 0.010$
Goodness of fit $\chi_{{\rm RVs}}^{2}/N_{{\rm RVs}}$ $2.3$
@@
Semi-major axis $a$ [au] $1.914\pm 0.018$
Photocenter semi-major axis $a_{0}$ [mas] $1.376\pm 0.016$
Eccentricity $e$ $0.6777\pm 0.0018$
Inclination $i$ [deg] $144.7\pm 0.5$
Periastron time $T_{p}$ [JD-2457389] $179.7\pm 2.3$
Ascending node angle $\Omega$ [deg] $70.5\pm 1.5$
Argument of periastron $\omega$ [deg] $334.36\pm 0.22$
Luminous star mass $M_{\star}$ [$M_{\odot}$] $0.781\pm 0.030$
@@
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $15.909\pm 0.034$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.1091\pm 0.0007$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $-264.202\pm 0.026$
@@
Orbital period $P_{\rm orb}$ [days] $827\pm 5$
Semi-major axis $a$ [au] $2.344\pm 0.031$
Photocenter semi-major axis $a_{0}$ [mas] $1.301\pm 0.017$
Eccentricity $e$ $0.357\pm 0.009$
Inclination $i$ [deg] $107.8\pm 1.0$
Periastron time $T_{p}$ [JD-2457389] $-47\pm 6$
Ascending node angle $\Omega$ [deg] $159.9\pm 1.4$
Argument of periastron $\omega$ [deg] $288.5\pm 2.4$
Luminous star mass $M_{\star}$ [$M_{\odot}$] $0.99\pm 0.05$
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.52\pm 0.07$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $19.0\pm 0.5$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.479\pm 0.030$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $56.8\pm 1.2$
Parallax $\varpi$ [mas] $0.916\pm 0.016$
Goodness of fit $\chi_{{\rm RVs}}^{2}/N_{{\rm RVs}}$ $1.4$
@@
Orbital period $P_{\rm orb}$ [days] $889.5\pm 0.7$
Semi-major axis $a$ [au] $2.405\pm 0.029$
Photocenter semi-major axis $a_{0}$ [mas] $5.923\pm 0.026$
Eccentricity $e$ $0.5840\pm 0.0035$
Inclination $i$ [deg] $125.29\pm 0.24$
Periastron time $T_{p}$ [JD-2457389] $-392.4\pm 0.9$
@@
Argument of periastron $\omega$ [deg] $243.03\pm 0.33$
@@
Neutron star mass $M_{2}$ [$M_{\odot}$] $1.396\pm 0.035$
RV semi-amplitude $K_{\star}$ [$\rm km\,s^{-1}$] $17.60\pm 0.12$
RV mass function $f\left(M_{2}\right)_{{\rm RVs}}$ [$\rm M_{\odot}$] $0.269\pm 0.005$
Center-of-mass RV $\gamma$ [$\rm km\,s^{-1}$] $-43.19\pm 0.11$
Parallax $\varpi$ [mas] $4.137\pm 0.016$
Goodness of fit $\chi_{{\rm RVs}}^{2}/N_{{\rm RVs}}$ $1.7$
@@
Orbital period $P_{\rm orb}$ [days] $1026.7\pm 3.3$
Semi-major axis $a$ [au] $2.776\pm 0.021$
Photocenter semi-major axis $a_{0}$ [mas] $2.701\pm 0.035$
Eccentricity $e$ $0.686\pm 0.012$